# Multimessenger Astronomy Beyond the Standard Model: New Window from Quantum Sensors

### v1

@@ VIII Conclusion and Outlook
In conclusion, we have laid the groundwork for exploring ULB signals in multimessenger searches involving astrophysical signals associated with GWs, photons, and neutrinos. Our analysis clarifies the role of screening and antiscreening in detection strategies, differentiates the behavior of linear and quadratic couplings, and underscores the potential of quantum sensor network observations in multimessenger astronomy. As experimental techniques and detection networks continue to improve, quantum sensors could provide valuable new insights into astrophysical events and the fundamental nature of beyond-SM quantum fields.
@@ A.0.1 Theory with $\phi^{6}$ potential
$d_{i}^{(2)}$ $\gtrsim M_{\text{pl}}^{2}\frac{\mathcal{R}^{-2}+\omega^{2}+4m^{2}}{32\pi\,\rho_{i}}~{}.$ (58)
@@ Appendix B Sensitivity for Additional Couplings
For some couplings (Figs. 12, 14, 15, 21, 22, 25, and 26), the best projections from proposed experiments are not sufficiently strong to overcome astrophysical constraints after rescaling the sensitivities for ULB bursts. This is especially relevant for sources at greater distances (e.g., $\mathcal{O}(10)$ Mpc) or for scalar linear couplings, where stringent constraints from EP tests Berge:2017ovy significantly limit the allowed parameter space. Multimessenger astronomy prospects with ultralight scalars are viable for linear couplings to photons and gluons for a Galactic source, and for quadratic couplings to photons, electrons and gluons. For ALPs, multimessenger astronomy prospects remain viable for the three nuclear couplings (to protons, neutrons, and nucleons via EDM), while the parameter space for electron and photon couplings is excluded due to tighter experimental and astrophysical constraints.
@@
Figure 11: Sensitivity to linear dilatonic scalar coupling to photons, $d_{e}^{(1)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{kpc}$ from Earth, releasing a total energy of $E_{\rm total}=10^{-2}M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The gray line and shaded region indicate the parameter space excluded by MICROSCOPE EP tests Berge:2017ovy . The black line shows the projected sensitivity of future experiments, based on $\eta=10^{-19}/5900$, consistent with a Thorium isomer nuclear clock expectation beeks2024fine , scaled for the scalar burst characteristics using Eq. (49). The black hatched and brown-shaded regions are excluded due to the scalar energy being less than its mass and the uncertainty principle, respectively. Dashed purple and red lines represent detection delays of 1 day and 1 year, respectively, relative to a signal traveling at the speed of light. The green shaded region denotes the viable parameter space for multimessenger astronomy with scalars, considering all constraints and the given burst benchmarks.
@@
Figure 12: Sensitivity to linear dilatonic scalar coupling to electrons, $d_{m_{e}}^{(1)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{kpc}$ from Earth, releasing a total energy of $E_{\rm total}=10^{-2}M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The gray line and shaded region indicate the parameter space excluded by MICROSCOPE EP tests Berge:2017ovy . The black line shows the projected sensitivity of future experiments, based on $\eta=10^{-17}$ from molecular spectra analysis of SrOH SrOH2021 , scaled for the scalar burst characteristics using Eq. (49). The black hatched and brown-shaded regions are excluded due to the scalar energy being less than its mass and the uncertainty principle, respectively. Dashed purple and red lines represent detection delays of 1 day and 1 year, respectively, relative to a signal traveling at the speed of light. For the given benchmarks, no viable parameter space exists for multimessenger astronomy with scalars.
@@
Figure 13: Sensitivity to linear dilatonic scalar coupling to gluons, $d_{g}^{(1)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{kpc}$ from Earth, releasing a total energy of $E_{\rm total}=10^{-2}M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The gray line and shaded region show the parameter space excluded by MICROSCOPE EP tests Berge:2017ovy . The black line represents the projected sensitivity of future experiments, based on $\eta=10^{-24}$, consistent with a Thorium isomer nuclear clock proposal 2012PhRvL.108l0802C ; FlaTh06 , scaled for the scalar burst characteristics using Eq. (49). The black hatched and brown shaded regions are excluded due to the scalar energy being less than its mass and the uncertainty principle, respectively. Dashed purple and red lines indicate detection delays of 1 day and 1 year, respectively, relative to a signal traveling at the speed of light. The green shaded region represents the viable parameter space for multimessenger astronomy with scalars, considering all constraints and the given benchmarks.
@@
Figure 14: Sensitivity to linear dilatonic scalar coupling to photons, $d_{e}^{(1)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{Mpc}$ from Earth, releasing a total energy of $E_{\rm total}=M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The gray region indicates parameter space excluded by MICROSCOPE EP tests Berge:2017ovy , and the black line represents the projected sensitivity of future experiments with $\eta=10^{-19}/5900$, based on a Thorium isomer nuclear clock proposal beeks2024fine , scaled using Eq. (49). The black hatched and brown regions are excluded due to the scalar energy being less than its mass and the uncertainty principle, respectively. Dashed purple and red lines correspond to detection delays of 1 day and 1 year relative to a light-speed signal. The green region denotes the viable parameter space for multimessenger astronomy with scalars, given all constraints and benchmarks.
@@
Figure 15: Sensitivity to linear dilatonic scalar coupling to electrons, $d_{m_{e}}^{(1)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{Mpc}$ from Earth, releasing a total energy of $E_{\rm total}=M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The gray region shows the parameter space excluded by MICROSCOPE EP tests Berge:2017ovy , and the black line indicates the projected sensitivity of future experiments with $\eta=10^{-17}$, based on molecular spectra analysis of SrOH SrOH2021 , scaled using Eq. (49). The black hatched and brown regions are excluded due to the scalar energy being less than its mass and the uncertainty principle, respectively. Dashed purple and red lines represent detection delays of 1 day and 1 year relative to a light-speed signal. For the given benchmarks, no viable parameter space exists for multimessenger astronomy with scalars.
@@
Figure 16: Sensitivity to linear dilatonic scalar coupling to gluons, $d_{g}^{(1)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{Mpc}$ from Earth, releasing a total energy of $E_{\rm total}=M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The gray region represents the parameter space excluded by MICROSCROPE EP tests Berge:2017ovy , and the black line indicates the projected sensitivity of future experiments with $\eta=10^{-24}$, based on a Thorium isomer nuclear clock proposal 2012PhRvL.108l0802C ; FlaTh06 , scaled using Eq. (49). The black hatched and brown regions are excluded due to the scalar energy being less than its mass and the uncertainty principle, respectively. Dashed purple and red lines correspond to detection delays of 1 day and 1 year relative to a light-speed signal. The green region indicates the viable parameter space for multimessenger astronomy with scalars, given all constraints and benchmarks.
@@
Figure 17: Sensitivity to quadratic dilatonic scalar coupling to electrons, $d_{m_{e}}^{(2)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{kpc}$ from Earth, releasing a total energy of $E_{\rm total}=10^{-2}M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The blue solid, dashed, and dotted lines represent critical screening due to the Earth, atmosphere, and experimental apparatus, respectively. The black line shows the projected sensitivity of future experiments with $\eta=10^{-17}$, based on molecular spectra analysis of SrOH SrOH2021 , scaled using Eq. (50). The black hatched and brown regions are excluded due to the scalar energy being less than its mass and the uncertainty principle, respectively. Dashed purple and red lines indicate detection delays of 1 day and 1 year relative to a light-speed signal. The green region represents the viable parameter space for multimessenger astronomy with scalars, considering all constraints and benchmarks.
@@
Figure 18: Sensitivity to quadratic dilatonic scalar coupling to gluons, $d_{g}^{(2)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{kpc}$ from Earth, releasing a total energy of $E_{\rm total}=10^{-2}M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The blue solid, dashed, and dotted lines indicate critical screening from the Earth, atmosphere, and experimental apparatus, respectively. The black line represents the projected sensitivity of future experiments with $\eta=10^{-24}$, based on a Thorium isomer nuclear clock proposal 2012PhRvL.108l0802C ; FlaTh06 , scaled using Eq. (50). The black hatched and brown regions are excluded by the requirements that scalar energy exceeds its mass and the uncertainty principle. Dashed purple and red lines show detection delays of 1 day and 1 year relative to a light-speed signal. The green region denotes the viable parameter space for multimessenger astronomy with scalars, considering all constraints and benchmarks.
@@
Figure 19: Sensitivity to quadratic dilatonic scalar coupling to electrons, $d_{m_{e}}^{(2)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{Mpc}$ from Earth, releasing a total energy of $E_{\rm total}=M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The blue solid, dashed, and dotted lines represent critical screening from the Earth, atmosphere, and experimental apparatus, respectively. The black line indicates the projected sensitivity of future experiments with $\eta=10^{-17}$, based on the SrOH molecular spectra analysis SrOH2021 , scaled using Eq. (50). The black hatched and brown regions are excluded by scalar energy limits and the uncertainty principle. Dashed purple and red lines show detection delays of 1 day and 1 year relative to a light-speed signal. The green region represents the viable parameter space for multimessenger astronomy with scalars, considering all constraints and benchmarks.
@@
Figure 20: Sensitivity to quadratic dilatonic scalar coupling to gluons, $d_{g}^{(2)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{Mpc}$ from Earth, releasing a total energy of $E_{\rm total}=M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The blue solid, dashed, and dotted lines indicate critical screening from the Earth, atmosphere, and experimental apparatus, respectively. The black line shows the projected sensitivity of future experiments with $\eta=10^{-24}$, based on a Thorium isomer nuclear clock proposal 2012PhRvL.108l0802C ; FlaTh06 , scaled using Eq. (50). The black hatched and brown regions are excluded by scalar energy constraints and the uncertainty principle. Dashed purple and red lines represent detection delays of 1 day and 1 year relative to a light-speed signal. The green region denotes viable parameter space for multimessenger astronomy with scalars, considering all constraints and benchmarks.
@@
Figure 21: Sensitivity to ALP derivative coupling to electrons, $g_{\phi e}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{kpc}$ from Earth, releasing a total energy of $E_{\rm total}=10^{-2}M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-20}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-15}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The gray region represents parameter space excluded by cooling in tip of the red-giant branch in $\omega$ Centauri from $\phi$ emission capozzi2020axion . The black line shows the projected sensitivity of future experiments with $g_{\rm DM}=2\times 10^{-14}~{}\mathrm{GeV}^{-1}$ and $t_{\rm int,DM}=1~{}\mathrm{year}$, based on proposed magnetometers chigusa2023light and torsion pendulums graham2018spin , scaled using Eq. (VI.2). The black hatched and brown regions are excluded due to ALP energy constraints and the uncertainty principle. Dashed purple and red lines indicate detection delays of 1 day and 1 year relative to a light-speed signal. For these benchmarks, no viable parameter space exists for multimessenger astronomy with ALPs.

### v2

@@ VIII Conclusion and Outlook
In conclusion, we have laid the groundwork for exploring ULB signals in multimessenger searches involving astrophysical signals associated with GWs, photons, and neutrinos. Our analysis clarifies the role of screening and antiscreening in detection strategies, differentiates the behavior of linear and quadratic couplings, and underscores the potential of quantum sensor network observations in multimessenger astronomy. Further, our analysis of screening also provides motivation and highlights the advantages of space-based detector systems, which avoid critical screening by the Earth and its atmosphere. As experimental techniques and detection networks continue to improve, quantum sensors could provide valuable new insights into astrophysical events and the fundamental nature of beyond-SM quantum fields.
@@ A.0.1 Theory with $\phi^{6}$ potential
$d_{i}^{(2)}$ $\gtrsim M_{\text{pl}}^{2}\frac{(2\mathcal{R})^{-2}+\omega^{2}+4m^{2}}{32\pi\,\rho_{i}}~{}.$ (58)
@@ Appendix B Sensitivity for Additional Couplings
For some couplings (Figs. 12, 14, 15, 25, and 26), the best projections from proposed experiments are not sufficiently strong to overcome astrophysical constraints after rescaling the sensitivities for ULB bursts. This is especially relevant for sources at greater distances (e.g., $\mathcal{O}(10)$ Mpc) or for scalar linear couplings, where stringent constraints from EP tests Berge:2017ovy significantly limit the allowed parameter space. Multimessenger astronomy prospects with ultralight scalars are viable for linear couplings to photons and gluons for a Galactic source, and for quadratic couplings to photons, electrons and gluons. For ALPs, multimessenger astronomy prospects remain viable for the three nuclear couplings (to protons, neutrons, and nucleons via EDM), while the parameter space for electron and photon couplings is excluded due to tighter experimental and astrophysical constraints.
@@
Figure 11: Sensitivity to linear dilatonic scalar coupling to photons, $d_{e}^{(1)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{kpc}$ from Earth, releasing a total energy of $E_{\rm total}=10^{-2}M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The gray line and shaded region indicate the parameter space excluded by MICROSCOPE EP tests Berge:2017ovy . The black line shows the projected sensitivity of future experiments, based on $\eta_{\mathrm{DM}}=10^{-19}/5900$, consistent with a Thorium isomer nuclear clock expectation beeks2024fine , scaled for the scalar burst characteristics using Eq. (49). The black hatched and brown-shaded regions are excluded due to the scalar energy being less than its mass and the uncertainty principle, respectively. Dashed purple and red lines represent detection delays of 1 day and 1 year, respectively, relative to a signal traveling at the speed of light. The green shaded region denotes the viable parameter space for multimessenger astronomy with scalars, considering all constraints and the given burst benchmarks.
@@
Figure 12: Sensitivity to linear dilatonic scalar coupling to electrons, $d_{m_{e}}^{(1)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{kpc}$ from Earth, releasing a total energy of $E_{\rm total}=10^{-2}M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The gray line and shaded region indicate the parameter space excluded by MICROSCOPE EP tests Berge:2017ovy . The black line shows the projected sensitivity of future experiments, based on $\eta_{\mathrm{DM}}=10^{-17}$ from molecular spectra analysis of SrOH SrOH2021 , scaled for the scalar burst characteristics using Eq. (49). The black hatched and brown-shaded regions are excluded due to the scalar energy being less than its mass and the uncertainty principle, respectively. Dashed purple and red lines represent detection delays of 1 day and 1 year, respectively, relative to a signal traveling at the speed of light. For the given benchmarks, no viable parameter space exists for multimessenger astronomy with scalars.
@@
Figure 13: Sensitivity to linear dilatonic scalar coupling to gluons, $d_{g}^{(1)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{kpc}$ from Earth, releasing a total energy of $E_{\rm total}=10^{-2}M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The gray line and shaded region show the parameter space excluded by MICROSCOPE EP tests Berge:2017ovy . The black line represents the projected sensitivity of future experiments, based on $\eta_{\mathrm{DM}}=10^{-24}$, consistent with a Thorium isomer nuclear clock proposal 2012PhRvL.108l0802C ; FlaTh06 , scaled for the scalar burst characteristics using Eq. (49). The black hatched and brown shaded regions are excluded due to the scalar energy being less than its mass and the uncertainty principle, respectively. Dashed purple and red lines indicate detection delays of 1 day and 1 year, respectively, relative to a signal traveling at the speed of light. The green shaded region represents the viable parameter space for multimessenger astronomy with scalars, considering all constraints and the given benchmarks.
@@
Figure 14: Sensitivity to linear dilatonic scalar coupling to photons, $d_{e}^{(1)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{Mpc}$ from Earth, releasing a total energy of $E_{\rm total}=M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The gray region indicates parameter space excluded by MICROSCOPE EP tests Berge:2017ovy , and the black line represents the projected sensitivity of future experiments with $\eta_{\mathrm{DM}}=10^{-19}/5900$, based on a Thorium isomer nuclear clock proposal beeks2024fine , scaled using Eq. (49). The black hatched and brown regions are excluded due to the scalar energy being less than its mass and the uncertainty principle, respectively. Dashed purple and red lines correspond to detection delays of 1 day and 1 year relative to a light-speed signal. For the given benchmarks, no viable parameter space exists for multimessenger astronomy with scalars.
@@
Figure 15: Sensitivity to linear dilatonic scalar coupling to electrons, $d_{m_{e}}^{(1)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{Mpc}$ from Earth, releasing a total energy of $E_{\rm total}=M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The gray region shows the parameter space excluded by MICROSCOPE EP tests Berge:2017ovy , and the black line indicates the projected sensitivity of future experiments with $\eta_{\mathrm{DM}}=10^{-17}$, based on molecular spectra analysis of SrOH SrOH2021 , scaled using Eq. (49). The black hatched and brown regions are excluded due to the scalar energy being less than its mass and the uncertainty principle, respectively. Dashed purple and red lines represent detection delays of 1 day and 1 year relative to a light-speed signal. For the given benchmarks, no viable parameter space exists for multimessenger astronomy with scalars.
@@
Figure 16: Sensitivity to linear dilatonic scalar coupling to gluons, $d_{g}^{(1)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{Mpc}$ from Earth, releasing a total energy of $E_{\rm total}=M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The gray region represents the parameter space excluded by MICROSCROPE EP tests Berge:2017ovy , and the black line indicates the projected sensitivity of future experiments with $\eta_{\mathrm{DM}}=10^{-24}$, based on a Thorium isomer nuclear clock proposal 2012PhRvL.108l0802C ; FlaTh06 , scaled using Eq. (49). The black hatched and brown regions are excluded due to the scalar energy being less than its mass and the uncertainty principle, respectively. Dashed purple and red lines correspond to detection delays of 1 day and 1 year relative to a light-speed signal. The green region indicates the viable parameter space for multimessenger astronomy with scalars, given all constraints and benchmarks.
@@
Figure 17: Sensitivity to quadratic dilatonic scalar coupling to electrons, $d_{m_{e}}^{(2)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{kpc}$ from Earth, releasing a total energy of $E_{\rm total}=10^{-2}M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The blue solid, dashed, and dotted lines represent critical screening due to the Earth, atmosphere, and experimental apparatus, respectively. The black line shows the projected sensitivity of future experiments with $\eta_{\mathrm{DM}}=10^{-17}$, based on molecular spectra analysis of SrOH SrOH2021 , scaled using Eq. (50). The black hatched and brown regions are excluded due to the scalar energy being less than its mass and the uncertainty principle, respectively. Dashed purple and red lines indicate detection delays of 1 day and 1 year relative to a light-speed signal. The green region represents the viable parameter space for multimessenger astronomy with scalars, considering all constraints and benchmarks.
@@
Figure 18: Sensitivity to quadratic dilatonic scalar coupling to gluons, $d_{g}^{(2)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{kpc}$ from Earth, releasing a total energy of $E_{\rm total}=10^{-2}M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The blue solid, dashed, and dotted lines indicate critical screening from the Earth, atmosphere, and experimental apparatus, respectively. The black line represents the projected sensitivity of future experiments with $\eta_{\mathrm{DM}}=10^{-24}$, based on a Thorium isomer nuclear clock proposal 2012PhRvL.108l0802C ; FlaTh06 , scaled using Eq. (50). The black hatched and brown regions are excluded by the requirements that scalar energy exceeds its mass and the uncertainty principle. Dashed purple and red lines show detection delays of 1 day and 1 year relative to a light-speed signal. The green region denotes the viable parameter space for multimessenger astronomy with scalars, considering all constraints and benchmarks.
@@
Figure 19: Sensitivity to quadratic dilatonic scalar coupling to electrons, $d_{m_{e}}^{(2)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{Mpc}$ from Earth, releasing a total energy of $E_{\rm total}=M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The blue solid, dashed, and dotted lines represent critical screening from the Earth, atmosphere, and experimental apparatus, respectively. The black line indicates the projected sensitivity of future experiments with $\eta_{\mathrm{DM}}=10^{-17}$, based on the SrOH molecular spectra analysis SrOH2021 , scaled using Eq. (50). The black hatched and brown regions are excluded by scalar energy limits and the uncertainty principle. Dashed purple and red lines show detection delays of 1 day and 1 year relative to a light-speed signal. The green region represents the viable parameter space for multimessenger astronomy with scalars, considering all constraints and benchmarks.
@@
Figure 20: Sensitivity to quadratic dilatonic scalar coupling to gluons, $d_{g}^{(2)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{Mpc}$ from Earth, releasing a total energy of $E_{\rm total}=M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The blue solid, dashed, and dotted lines indicate critical screening from the Earth, atmosphere, and experimental apparatus, respectively. The black line shows the projected sensitivity of future experiments with $\eta_{\mathrm{DM}}=10^{-24}$, based on a Thorium isomer nuclear clock proposal 2012PhRvL.108l0802C ; FlaTh06 , scaled using Eq. (50). The black hatched and brown regions are excluded by scalar energy constraints and the uncertainty principle. Dashed purple and red lines represent detection delays of 1 day and 1 year relative to a light-speed signal. The green region denotes viable parameter space for multimessenger astronomy with scalars, considering all constraints and benchmarks.
@@
Figure 21: Sensitivity to ALP derivative coupling to electrons, $g_{\phi e}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{kpc}$ from Earth, releasing a total energy of $E_{\rm total}=10^{-2}M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-20}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-15}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The gray region represents parameter space excluded by cooling in tip of the red-giant branch in $\omega$ Centauri from $\phi$ emission capozzi2020axion . The black line shows the projected sensitivity of future experiments with $g_{\rm DM}=2\times 10^{-14}~{}\mathrm{GeV}^{-1}$ and $t_{\rm int,DM}=1~{}\mathrm{year}$, based on proposed magnetometers chigusa2023light and torsion pendulums graham2018spin , scaled using Eq. (VI.2). The black hatched and brown regions are excluded due to ALP energy constraints and the uncertainty principle. Dashed purple and red lines indicate detection delays of 1 day and 1 year relative to a light-speed signal. The green region denotes viable parameter space for multimessenger astronomy with scalars, considering all constraints and benchmarks.